# thanks: Corresponding author

### v1

@@ I Introduction
The study of cosmology has entered the era of precision cosmology Spergel et al. (2003); Bennett et al. (2003). Five of the six basic parameters of the $\Lambda$CDM model have been precisely constrained by the Planck cosmic microwave background (CMB) data with constraint precisions better than 1% Aghanim et al. (2020). However, as the precision of cosmological parameter measurements improves, some puzzling tensins have come to light. The Hubble constant ($H_{0}$) is a crucial cosmological parameter describing the universe’s current expansion rate. Its values inferred from the Planck CMB observation (assuming the $\Lambda$CDM model) Aghanim et al. (2020) and the model-independent local-universe distance ladder measurement Riess et al. (2022) are in more than $5\sigma$ tension, which has become a critical crisis in cosmology Verde et al. (2019); Riess (2019) (see Refs. Li et al. (2013); Zhang et al. (2015, 2014a, 2014b); Zhao et al. (2017); Guo and Zhang (2017); Guo et al. (2019, 2020); Yang et al. (2018); Verde et al. (2019); Riess (2019); Di Valentino et al. (2020a, b); Liu et al. (2020); Zhang and Huang (2020); Ding et al. (2020); Cai (2020); Di Valentino et al. (2021); Abdalla et al. (2022); Cai et al. (2021); Li and Zhang (2020); Wang et al. (2022a); Vagnozzi et al. (2022); Vagnozzi (2021, 2020); Feng et al. (2020a); Lin et al. (2020); Hryczuk and Jodłowski (2020); Gao et al. (2021); Zhao et al. (2022); Gao et al. (2024) for related discussions). While the Hubble tension indicates the potential for new physics beyond the $\Lambda$CDM model, there remains a lack of consensus on a valid extended cosmological model to address this tension. Additionally, it is imperative to explore new cosmological probes capable of independently measuring $H_{0}$.
@@
Compact binary coalescence (CBC) events, a category of gravitational wave (GW) sources, have been discovered by the LIGO-Virgo-KAGRA collaboration in recent years, with approximately a hundred events detected. The luminosity distances of CBC events can be determined through the analysis of GW waveforms, and these events are commonly referred to as “standard sirens” Schutz (1986); Holz and Hughes (2005). Standard sirens can provide absolute distance measurements, avoiding poorly understood calibration processes inherent in the distance ladder method. If the redshifts of GW sources are also determined, the distance-redshift relation could be established to study the expansion history of the universe (see Refs. Dalal et al. (2006); Cutler and Holz (2009); Zhao et al. (2011); Cai and Yang (2017, 2018); Cai et al. (2018a); Zhao et al. (2018); Wang et al. (2018); Zhang et al. (2019a, 2020); Li et al. (2020); Cai et al. (2018b); Chang et al. (2019); Du et al. (2019); He (2019); Yang et al. (2019, 2020); Zhang (2019); Bachega et al. (2020); Zhang et al. (2019b); Borhanian et al. (2020); Chen (2020); Jin et al. (2020); Chen et al. (2021); Mitra et al. (2021); Zhao et al. (2020a); Jin et al. (2021, 2022, 2023a, 2023b, 2023c, 2024); Qi et al. (2021); Ye and Fishbach (2021); Cao et al. (2022); de Souza et al. (2022); Wang et al. (2022b); Colgáin (2022); Dhani et al. (2022); Wang et al. (2022c); Zhu et al. (2022); Califano et al. (2023); Hou et al. (2023); Wu et al. (2023); Han et al. (2023); Zhang et al. (2023); Li et al. (2024, 2024); Yu et al. (2024) for related discussions). The binary neutron star coalescence event, GW170817, gives $H_{0}=70^{+12}_{-8}~{}\rm~{}km~{}s^{-1}~{}Mpc^{-1}$ with a precision of 14% Abbott et al. (2017a). Its redshift was determined by directly identifying the electromagnetic (EM) signals produced when the binary merged Abbott et al. (2017b, c). Such standard sirens are known as bright sirens, with only one case detected to date. For the standard sirens without EM counterparts (usually referred to as dark sirens), we can cross-match them with galaxy catalogs and statistically infer their redshifts (see Refs. Del Pozzo (2012); Chen et al. (2018); Nair et al. (2018); Fishbach et al. (2019); Gray et al. (2020); Soares-Santos et al. (2019); Abbott et al. (2021); Yu et al. (2020); Palmese et al. (2020, 2023); Abbott et al. (2023); Leandro et al. (2022); Finke et al. (2021); Gair et al. (2023); Song et al. (2024); Yang et al. (2022); Muttoni et al. (2023) for related works). Recently, using 46 dark sirens with signal-to-noise ratios (SNRs) over 11 from GWTC-3 and the GLADE+ $K$-band galaxy catalog Dálya et al. (2018, 2022), researchers achieved a $\sim$19constraint on $H_{0}$. Combined with the bright siren GW170817, the constraint precision of $H_{0}$ improves to $\sim$10Abbott et al. (2023).
@@
The frequencies of GWs generated by the GW sources with varying masses are hugely different, which requires GW detectors sensitive to diverse frequency bands. The current LIGO-Virgo-KAGRA detector network Somiya (2012); Aasi et al. (2015); Acernese et al. (2015) can detect only stellar-mass CBCs at around hundreds of hertz. In the future, precise measurements of GWs in various bands will become possible. In the frequency band from a few hertz to several thousand hertz, the third-generation ground-based GW detectors, e.g., the Cosmic Explorer Abbott et al. (2017d) and the Einstein Telescope Punturo et al. (2010), will significantly enhance the detection capabilities for stellar-mass CBCs. In the decihertz band, the Decihertz Interferometer Gravitational-Wave Observatory Seto et al. (2001); Kawamura et al. (2019); Dong et al. (2024) and lunar-based detectors Harms et al. (2021); Jani and Loeb (2021) could detect intermediate-mass black hole binaries and binary white dwarfs. In the millihertz band, GWs from massive black hole binaries and extreme mass ratio inspirals could be detected by space-borne GW detectors such as the Laser Interferometer Space Antenna Amaro-Seoane et al. (2017); Robson et al. (2019); Auclair et al. (2023), Taiji Sirunyan et al. (2018); Ruan et al. (2020); Wu (2018); Zhao et al. (2020b), and TianQin Luo et al. (2016); Wang et al. (2019, 2020); Luo et al. (2020); Milyukov (2020); Mei et al. (2021). In the nanohertz (nHz) band, the pulsar timing arrays (PTAs) could detect continuous GWs from the inspiraling supermassive black hole binaries (SMBHBs) resulting from galaxy mergers.
@@
The pulsars constructing PTAs are millisecond pulsars (MSPs) and have highly stable rotation periods. Their radio signals can be periodically detected by radio telescopes on the Earth with extremely high precision. GWs passing between the MSPs and the Earth will alter the proper distances between them. This leads to a discrepancy between the actual and expected arriving times of radio signals, which is known as the timing residuals. By analyzing the time residuals of PTAs, we can extract information about nHz GWs. Recently, North American Nanohertz Observatory for Gravitational Waves Agazie et al. (2023a, b), European PTA + Indian PTA Antoniadis et al. (2023a, b), Parkes PTA Zic et al. (2023); Reardon et al. (2023), and Chinese PTA Xu et al. (2023) have reported substantial evidence for the existence of a stochastic GW background (SGWB) Hellings and Downs (1983). Such evidence boosts confidence in making more advances in detecting nHz GWs. For instance, detecting individual SMBHBs has become one of the next important missions of PTA observations, which has been widely discussed in Refs. Sesana et al. (2009); Corbin and Cornish (2010); Sesana and Vecchio (2010); Babak and Sesana (2012); Lee et al. (2011); Ellis et al. (2012); Ellis (2013); Zhu et al. (2015); Wang and Mohanty (2017); Feng et al. (2020b). Although extracting an individual GW signal from the superimposed GW background is challenging, this goal may be achievable with the addition of more advanced radio telescopes, especially in the era of the Square Kilometre Array (SKA).
@@
Recently, Refs. Yan et al. (2020); Wang et al. (2022d) indicated that utilizing SMBHBs as standard sirens could help to improve the measurement precisions of cosmological parameters. The authors of Ref. Yan et al. (2020) pointed out that the EM counterparts of SMBHBs can be observed during its inspiraling stage and forecast the constraint on the equation of state of dark energy with 154 SMBHB bright sirens observed by SKA-era PTAs. The authors of Ref. Wang et al. (2022d) investigated the dominant factors affecting the detection capability of PTAs for SMBHBs and found that the root mean square (RMS) of white noise, $\sigma_{t}$, is more dominant than the number of pulsars in affecting the detections of SMBHBs. They also proposed that detecting nHz GWs from individual SMBHBs has significant implications for exploring the properties of dark energy and precisely determining the Hubble constant.
@@
In the aforementioned papers, the SMBHB standard siren mock data are simulated based on the galactic major merger model proposed by Mingarelli et al. Mingarelli et al. (2017) (hereafter the M17 model). Recently, Casey-Clyde et al. Casey-Clyde et al. (2022) indicated that most major merger models predicted characteristic amplitudes of SGWB lower than that derived from the NANOGrav 12.5-yr data. In this context, they proposed a quasar-based SMBHB population model (hereafter the C21 model), considering that galaxy mergers could trigger quasars, and SMBHBs arising from galaxy mergers might be linked to quasar populations. This model can self-consistently predict the local number density of SMBHBs given the amplitude of the SGWB.
@@
Based on the C21 model and the common-process signal in the NANOGrav 12.5-yr dataset, the number of SMBHBs in the local universe is roughly five times higher than that predicted by the M17 model. It is evident that there is a significant difference in the number of SMBHBs predicted by these two models. Notably, the number of SMBHB dark sirens significantly influences the constraints on cosmological parameters, especially when the localization of dark sirens is poor Abbott et al. (2023). Nevertheless, previous studies have not forecast the capability of SMBHB dark sirens of constraining cosmological parameters based on the C21 model. Consequently, a crucial question arises: Based on the C21 model that is more consistent with SGWB observations, what role can SMBHB dark sirens actually play in measuring cosmological parameters?
@@
To answer this question, we simulate the mock SMBHB dark siren data based on the C21 quasar-based model and the SGWB signal inferred from the NANOGrav 15-yr dataset Agazie et al. (2023a). We analyze the ability of SKA-era PTAs to detect the local-universe SMBHBs based on various PTA detection cases, and then combine the mock dark siren data with the 2 Micron All Sky Survey (2MASS) Extended Source Catalog Jarrett et al. (2000) to perform cosmological parameter estimations. In this work, we only focus on the ability of SMBHB dark sirens to measure $H_{0}$. This is because our aim is to investigate how SMBHB dark sirens can contribute to resolving the Hubble tension, and we primarily mock SMBHB dark sirens in the local universe, where the distance-redshift relationship is only sensitive to $H_{0}$. Therefore, we use the SMBHBs dark sirens to infer $H_{0}$ and fix all other cosmological parameters to the Planck 2018 values.
@@
This paper is organized as follows. Section II introduces the methodology used in this work. In Section III, we report our constraint results and make relevant discussions. The conclusion is given in Section IV. Unless otherwise stated, we adopt the system of units in which $G=c=1$ throughout this work.
II METHODOLOGY
II.1 Subcatalog of 2MASS galaxy catalog
@@
Massive galaxies are hypothesized to have a higher probability of hosting SMBHBs Rosado and Sesana (2014). To select massive galaxies in the local universe, we first estimate their $K$-band absolute magnitudes by $M_{K}=m_{K}-5\times\log_{10}(d_{\rm L})-25-0.11\times A_{v}$, where $d_{\rm L}$ represents the luminosity distance to the galaxy. Ref. Arzoumanian et al. (2021) considered the impact of peculiar motions on redshift measurements, allowing for precise calculations of distances for all galaxies in the 2MASS Redshift Survey (2MRS) catalog Huchra et al. (2012). The apparent magnitude $m_{K}$ and the extragalactic extinction $A_{v}$ can be obtained from the 2MRS catalog. We implement a $K$-band cutoff with $M_{K}\leq-25$ for selecting galaxies with stellar mass $M_{\star}\geq 10^{11}M_{\odot}$. Based on this criterion, we select 13172 galaxies from the 2MRS catalog as potential host galaxies for SMBHBs. These candidate host galaxies constitute a new catalog, referred to as the 2MASS catalog in the following text.
@@
$P_{\text{com}}(\mathcal{S})\equiv\frac{N_{\text{cat}}(\mathcal{S})}{\bar{n}_{\mathrm{gal}}V_{\rm c}(\mathcal{S})},$ (1)
@@
where $\mathcal{S}=\mathcal{S}(z,{\hat{\Omega}};\Delta z,\Delta\Omega)$ represents a volume that is a portion of a cone. This cone is defined by the central redshift $z$ with a range from $z-\Delta z$ to $z+\Delta z$, and the central angle ${\hat{\Omega}}$ with a solid angle $\Delta\Omega$. $\bar{n}_{\text{gal}}$ represents the density of galaxies per comoving volume. $V_{\rm c}$ is the comoving volume of $\mathcal{S}$, and $N_{\text{cat}}$ is the number of galaxies in the catalog within $\mathcal{S}$. We assume that the completeness of the 2MASS catalog is isotropic and count the number of galaxies in each redshift bin. The redshift range from 0 to 0.1 is equally divided into 30 bins. Additionally, we assume that galaxies are uniformly distributed within the comoving volume and consider the 2MASS catalog complete at $z<0.04$. Therefore, we can derive the completeness of the 2MASS catalog as a function of redshift using Eq. (1), which is shown in Fig. 1.
@@
Figure 1: The red solid lines represent the subcatalog completeness defined in Eq. (1) as a function of redshift. The histogram shows the number of galaxies within each redshift bin in the subcatalog. The blue dashed line represents the number of galaxies expected within each redshift bin assuming a uniform distribution of galaxies in comoving volume.
II.2 Population models of SMBHBs

### v2

# Nanohertz gravitational waves from a quasar-based supermassive black hole binary population model as dark sirens

@@ 1 Introduction
The study of cosmology has entered the era of precision cosmology [1, 2]. Five of the six basic parameters of the $\Lambda$CDM model have been precisely constrained by the Planck cosmic microwave background (CMB) data with constraint precisions better than 1% [3]. However, as the precision of cosmological parameter measurements improves, some puzzling tensions have come to light. The Hubble constant ($H_{0}$) is a crucial cosmological parameter describing the universe’s current expansion rate. Its values inferred from the Planck CMB observation assuming the $\Lambda$CDM model [3] and the model-independent local-universe distance ladder measurement [4] are in more than $5\sigma$ tension, which has become a critical crisis in cosmology [5, 6] and has been widely discussed (see, e.g. [7, 8, 9, 10, 11, 12, 13, 14, 5, 6, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]). While the Hubble tension indicates the potential for new physics beyond the $\Lambda$CDM model, there remains a lack of consensus on a valid extended cosmological model to address this tension. Additionally, it is imperative to explore new cosmological probes capable of independently measuring $H_{0}$.
@@
Compact binary coalescence (CBC) events, a category of gravitational wave (GW) sources, have been discovered by the LIGO-Virgo-KAGRA collaboration in recent years, with approximately a hundred events detected. The luminosity distances of CBC events can be determined through the analysis of GW waveforms, and these events are commonly referred to as “standard sirens” [36, 37]. Standard sirens can provide absolute distance measurements, avoiding poorly understood calibration processes inherent in the distance ladder method. If the redshifts of GW sources are also determined, the distance-redshift relation could be established to study the expansion history of the universe (see, e.g. [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84]). The binary neutron star coalescence event, GW170817, gives $H_{0}=70^{+12}_{-8}~{}\rm~{}km~{}s^{-1}~{}Mpc^{-1}$ with a precision of 14% [85]. Its redshift was determined by directly identifying the electromagnetic (EM) signals produced when the binary merged [86, 87]. Such standard sirens are known as bright sirens, with only one case detected to date.
For the standard sirens without EM counterparts (usually referred to as dark sirens), we can cross-match them with galaxy catalogs and statistically infer their redshifts (see, e.g. [88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104]). Recently, using 46 dark sirens with signal-to-noise ratios (SNRs) over 11 from GWTC-3 and the GLADE+ $K$-band galaxy catalog [105, 106], researchers achieved a $\sim$19constraint on $H_{0}$. Combined with the bright siren GW170817, the constraint precision of $H_{0}$ improves to $\sim$10[98].
@@
The frequencies of GWs generated by the GW sources with varying masses are hugely different, which requires GW detectors sensitive to diverse frequency bands. The current LIGO-Virgo-KAGRA detector network [107, 108, 109] can detect only stellar-mass CBCs at around hundreds of hertz. In the future, precise measurements of GWs in various bands will become possible. In the frequency band from a few hertz to several thousand hertz, the third-generation ground-based GW detectors, e.g., the Cosmic Explorer [110] and the Einstein Telescope [111], will significantly enhance the detection capabilities for stellar-mass CBCs. In the decihertz band, the Decihertz Interferometer Gravitational-Wave Observatory [112, 113, 114] and lunar-based detectors [115, 116] could detect intermediate-mass black hole binaries and binary white dwarfs. In the millihertz band, GWs from massive black hole binaries and extreme mass ratio inspirals could be detected by space-borne GW detectors such as the Laser Interferometer Space Antenna [117, 118, 119], Taiji [120, 121, 122, 123], and TianQin [124, 125, 126, 127, 128, 129]. In the nanohertz (nHz) band, the pulsar timing arrays (PTAs) could detect continuous GWs from the inspiraling supermassive black hole binaries (SMBHBs) resulting from galaxy mergers.
@@
The pulsars constructing PTAs are millisecond pulsars (MSPs) and have highly stable rotation periods. Their radio signals can be periodically detected by radio telescopes on the Earth with extremely high precision. GWs passing between the MSPs and the Earth will alter the proper distances between them. This leads to a discrepancy between the actual and expected arriving times of radio signals, which is known as the timing residuals. By analyzing the time residuals of PTAs, we can extract information about nHz GWs. Recently, North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [130, 131], European PTA (EPTA) + Indian PTA (InPTA) [132, 133], Parkes PTA (PPTA) [134, 135], and Chinese PTA (CPTA) [136] have reported substantial evidence for the existence of a stochastic GW background (SGWB) [137]. Such evidence boosts confidence in making more advances in detecting nHz GWs. For instance, detecting individual SMBHBs has become one of the next important missions of PTA observations [138, 139, 140, 141, 142, 143, 144, 145, 146, 147]. Although extracting an individual GW signal from the superimposed GW background is challenging, this goal may be achievable with the addition of more advanced radio telescopes, especially in the era of the Square Kilometre Array (SKA).
@@
Recently, Yan et al. [148] and Wang et al. [149] indicated that utilizing SMBHBs as standard sirens could help to improve the measurement precisions of cosmological parameters. Yan et al. [148] pointed out that the EM counterparts of SMBHBs can be observed during its inspiraling stage and forecast the constraint on the equation of state of dark energy with 154 SMBHB bright sirens observed by SKA-era PTAs. Wang et al. [149] investigated the dominant factors affecting the detection capability of PTAs for SMBHBs and found that the root mean square (RMS) of white noise, $\sigma_{t}$, is more dominant than the number of MSPs in affecting the detections of SMBHBs. They also proposed that detecting nHz GWs from individual SMBHBs has significant implications for exploring the properties of dark energy and precisely determining the Hubble constant.
@@
In the aforementioned papers, the SMBHB standard siren mock data are simulated based on the galactic major merger model proposed by Mingarelli et al. [150] (hereafter the M17 model). Recently, Casey-Clyde et al. [151] indicated that most major merger models predicted characteristic amplitudes of SGWB lower than that derived from the NANOGrav 12.5-yr data. In this context, they proposed a quasar-based SMBHB population model (hereafter the C21 model), considering that galaxy mergers could trigger quasars, and SMBHBs arising from galaxy mergers might be linked to quasar populations. This model can self-consistently predict the local number density of SMBHBs given the amplitude of the SGWB.
@@
Based on the C21 model and the common-process signal in the NANOGrav 12.5-yr dataset, the number of SMBHBs in the local universe is roughly five times higher than that predicted by the M17 model. It is evident that there is a significant difference in the number of SMBHBs predicted by these two models. Notably, the number of SMBHB dark sirens significantly influences the constraints on cosmological parameters, especially when the localization of dark sirens is poor [98]. Nevertheless, previous studies have not forecast the capability of SMBHB dark sirens of constraining cosmological parameters based on the C21 model. Consequently, a crucial question arises: Based on the C21 model that is more consistent with SGWB observations, what role can SMBHB dark sirens actually play in measuring cosmological parameters?
@@
To answer this question, we simulate the mock SMBHB dark siren data based on the C21 quasar-based model and the SGWB signal inferred from the NANOGrav 15-yr dataset [130]. It needs to be emphasized that we employ the latest NANOGrav 15-yr dataset rather than the 12.5-yr dataset to ensure that our SMBHB population estimate is consistent with the most up-to-date PTA observations. We analyze the ability of SKA-era PTAs to detect the local-universe SMBHBs based on various PTA detection cases, and then combine the mock dark siren data with the 2 Micron All Sky Survey (2MASS) Extended Source Catalog [152] to perform cosmological parameter estimations. In this work, we focus only on the ability of SMBHB dark sirens to measure $H_{0}$. This is because our aim is to investigate how SMBHB dark sirens can contribute to resolving the Hubble tension, and we primarily mock SMBHB dark sirens in the local universe, where the distance-redshift relationship is sensitive only to $H_{0}$. Therefore, we use the SMBHBs dark sirens to infer $H_{0}$ and fix all other cosmological parameters to the Planck 2018 values.
@@
This paper is organized as follows. Section 2 introduces the methodology used in this work. In Section 3, we report our constraint results and make relevant discussions. The conclusion is given in Section 4. Unless otherwise stated, we adopt the system of units in which $G=c=1$ throughout this work.
2 Methodology
2.1 Subcatalog of 2MASS galaxy catalog
@@
Massive galaxies are hypothesized to have a higher probability of hosting SMBHBs [153]. To select massive galaxies in the local universe, we first estimate their $K$-band absolute magnitudes by $M_{K}=m_{K}-5\times\log_{10}(d_{\rm L})-25-0.11\times A_{v}$, where $d_{\rm L}$ represents the luminosity distance to the galaxy. The NANOGrav Collaboration [154] considered the impact of peculiar velocities on redshift measurements, allowing for precisely calculating the distances of all galaxies in the 2MASS Redshift Survey (2MRS) catalog [155]. The apparent magnitude $m_{K}$ and the extragalactic extinction $A_{v}$ can be obtained from the 2MRS catalog. We implement a $K$-band cutoff with $M_{K}\leq-25$ for selecting galaxies with stellar mass $M_{\star}\geq 10^{11}M_{\odot}$. Based on this criterion, we select 13172 galaxies from the 2MRS catalog as potential host galaxies for SMBHBs. These candidate host galaxies constitute a new catalog, referred to as the 2MASS catalog in the following text.
@@
$P_{\text{com}}(\mathcal{S})\equiv\frac{N_{\text{cat}}(\mathcal{S})}{\bar{n}_{\mathrm{gal}}V_{\rm c}(\mathcal{S})},$ (2.1)
@@
where $\mathcal{S}=\mathcal{S}(z,{\hat{\Omega}};\Delta z,\Delta\Omega)$ represents a volume that is a portion of a cone. This cone is defined by the central redshift $z$ with a range from $z-\Delta z$ to $z+\Delta z$, and the central angle ${\hat{\Omega}}$ with a solid angle $\Delta\Omega$. $\bar{n}_{\text{gal}}$ represents the density of galaxies per comoving volume. $V_{\rm c}$ is the comoving volume of $\mathcal{S}$, and $N_{\text{cat}}$ is the number of galaxies in the catalog within $\mathcal{S}$. We assume that the completeness of the 2MASS catalog is isotropic and count the number of galaxies in each redshift bin. The redshift range from 0 to 0.1 is equally divided into 30 bins. Additionally, we assume that galaxies are uniformly distributed within the comoving volume and consider the 2MASS catalog complete at $z<0.04$. Therefore, we can derive the completeness of the 2MASS catalog as a function of redshift using Eq. (2.1), which is shown in figure 1.
@@
Figure 1: The red solid lines represent the subcatalog completeness defined in Eq. (2.1) as a function of redshift. The histogram shows the number of galaxies within each redshift bin in the subcatalog. The blue dashed line represents the number of galaxies expected within each redshift bin assuming a uniform distribution of galaxies in comoving volume.
2.2 Population models of SMBHBs